Question:

A particle moves such that its vector is given by ๐‘Ÿ

ย = cos ๐œ”๐‘ก๐‘– + sin ๐œ”๐‘ก๐‘— where ๐œ” is a constant. Show that:

(i)

Velocity, ๐‘ฃย of the particle is perpendicular toย r

(ii)

Acceleration,ย ๐‘Žย is directed towards the origin and has magnitude proportional to theย 

distance from the origin

(iii) ๐‘Ÿ ๐‘‹ ๐‘ฃis a constant vector



1
Expert's answer
2021-08-30T16:09:38-0400

(i)


"v=\dfrac{dr}{dt}=-\omega\sin \omega ti+\omega\cos \omega tj"

"r\cdot v=\cos \omega t(-\omega\sin \omega t)+\sin \omega t(\omega \cos \omega t)=0"

Hence "r \perp v."


(ii)


"a=\dfrac{d^2r}{dt^2}=\dfrac{dv}{dt}=-\omega^2\cos \omega ti-\omega^2\sin \omega tj"

"=-\omega^2(\cos \omega ti+\sin\omega tj)=-\omega^2r"

Hence ๐‘Žย is directed towards the origin and has magnitude proportional to theย  distance from the origin


"a=(-\omega^2)r"

(iii)


"r\times v=\begin{vmatrix}
 i & j & k \\
 \cos\omega t & \sin\omega t & 0\\
 -\omega \sin\omega t & \omega\cos\omega t & 0\\
\end{vmatrix}"


"=k\begin{vmatrix}
 \cos\omega t & \sin\omega t \\
 -\omega \sin\omega t & \omega \cos\omega t 
\end{vmatrix}"

"=(\omega\cos^2\omega t+\omega\sin^2\omega t)k"

"=\omega k"

Therefore "r\times v" is a constant vector ("\omega=const" ).